# Observe Then Act: Asynchronous Active Vision-Action Model for Robotic Manipulation

### v1

@@ VI Result
TABLE V: Ablation over voxel alignment,demo augmentation and task-agnostic auxiliary rewards
Models Success Rate Episode Length Overtime Rate Planning Failed Rate
Ours(w/o align) 0.34±0.41 7.61±2.94 0.14±0.12 0.53±0.39
Ours(w/o aug) 0.31±0.37 7.99±2.44 0.15±0.17 0.55±0.37
Ours(w/o aux) 0.70±0.19 5.52±1.51 0.15±0.10 0.16±0.12
Ours (w/o align+aug+aux) 0.22±0.34 8.46±2.41 0.22±0.22 0.57±0.34
Ours 0.76±0.15 4.88±1.16 0.10±0.06 0.15±0.11
Models First Interaction Non-interaction AIIG ASSIG
Ours(w/o align) 4.33±2.80 5.81±3.30 0.046±0.05 0.082±0.07
Ours(w/o aug) 4.49±3.34 5.74±3.58 0.014±0.05 0.020±0.03
Ours(w/o aux) 1.85±0.96 2.38±1.25 0.067±0.08 0.028±0.02
Ours (w/o align+aug+aux) 5.74±3.37 6.72±3.28 0.036±0.07 0.083±0.06
Ours 1.64±0.61 2.15±1.03 0.068±0.09 0.035±0.03
VII Conclusion and Feature Work
@@
In this paper, we investigate the problem of robotic manipulation under limited observation and present a task-driven asynchronous active vision-action model. Our model first infers the optimal observation viewpoint accroding to the task goal and then determines the corresponding manipulation action based on the updated viewpoint’s obsevation. To enhance performance and stability, our model integrates three strategies: voxel alignment for spatial consistency, demonstration augmentation for better learning, and auxiliary rewards for real-time feedback. Together, these strategies help the model handle occlusions and improve sensor-motor coordination. Experimental results show that, compared to passive observation-based model, our model significantly improves manipulation performance through actively observation. It further learns sensor-motor coordination through interaction with the environment, rather than being restricted to a limited set of demonstration trajectories. Despite these promising results, the model’s performance can still be affected by complex occlusions and dynamic environments. Future research could explore incorporating real-time visual feedback to better handle dynamic changes in the environment.

### v2

@@ VI Result
TABLE IV: Ablation over voxel alignment,demo augmentation and task-agnostic auxiliary rewards.
Ablation Setting Result Proportion Interactivity Sensing
align aug aux SR EF TO EL FI NI AIIG ASSIG
0.34 0.53 0.14 7.61 4.33 5.81 0.046 0.082
0.31 0.55 0.15 7.99 4.49 5.74 0.014 0.020
0.70 0.16 0.15 5.52 1.85 2.38 0.067 0.028
0.22 0.57 0.22 8.46 5.74 6.72 0.036 0.083
0.76 0.15 0.09 4.88 1.64 2.15 0.068 0.035
VII Conclusion and Future Work
@@
In this paper, we investigate the problem of robotic manipulation under limited observation and present a task-driven asynchronous active vision-action model. Our model first infers the optimal viewpoint based on the task goal, then selects the manipulation action from the updated observation. To enhance performance and training stability, we integrate voxel alignment for spatial consistency, demonstration augmentation for enhanced learning, and auxiliary rewards for real-time training guidance. These strategies enable the model to handle occlusions and improve sensor-motor coordination. Experimental results show that our model significantly outperforms passive observation-based models by actively refining viewpoints and learning sensor-motor coordination through interaction with the environment. However, complex occlusions and dynamic environments remain challenging. Future work could explore real-time visual feedback to better adapt to environmental changes.